In 2017, labor productivity in Quebec (in Canadian dollars, at purchasing power parity) amounted to $ 60.15. For the Canadian average, we are talking about a GDP per hour worked of $ 66.96.

Our productivity is much lower compared to that of the United States ($ 82.03), Germany ($ 87.38) or Belgium ($ 93.43). Quebec is not at the bottom of the rankings, however: they are South Korea ($ 45.77) and Japan ($ 59.46).

In 2009, labor productivity in Quebec was $ 48.56. What is the productivity growth rate in Quebec between 2009 and 2017?
1
Expert's answer
2021-06-07T12:39:05-0400

Productivity rate

Refers to the work output divided by the number of inputs such as labour or capital.


"Productivity = \frac{Output}{inputs}"


In the above case the productivity of an entire county("\frac{output}{Inputs})" is already calculated.


Considering the changes in time, to find the growth rate, the difference between 2009 and 2017 is first computed then divided by time ( number of years)


Let's do it;

"\frac{60.15 - 48.56}{ 8}"


Productivity growth rate "=1.45"